OpenStudy (anonymous):

Write the equation of the line in point-slope form for the line through the given point (4, -6) and with m = 3/5

OpenStudy (anonymous):

Okay, one minute. :)

OpenStudy (anonymous):

Point slope form is: y - y1 = m(x - x1) y - (-6) = 3/5(x - (4)) y + 6 = 3/5(x - 4) That is your answer. c: